A penny is 1.55 mm thick.
 Imagine that you stack pennies on top of each other every day. Each day, you double the  amount of pennies in the stack. This means that on day 0, there is 1 penny. On day 1, there  are 2 pennies. On day 2, there are 4 pennies, etc. What exponential function is being modeled  here?

Answer by jorel555(1290)

If n equals the number of days, then:
1.55*2^n=1000
2^n=645.16
2^9=512, 2^10=1024
Your stack will be over a meter tall after the 9th day!!!!!!!!!!